# The electrostatic charge on a falling water drop

### v1

@@ Abstract
Fluid triboelectrification, also known as flow electrification, remains an under-explored yet ubiquitous phenomenon with potential applications in various fields, from material science to planetary evolution. Building upon previous efforts to position water within the triboelectric series, we investigate the charge on individual, millimetric water drops falling through air. Our experiments measured the charge and mass of each drop using a Faraday cup mounted on a mass balance, and connected to an electrometer. For pure water in a glass syringe with a grounded metal tip, we find the charge per drop ($\Delta q/\Delta m$) was approximately -5 pC/g to -1 pC/g. This was independent of the release height of the drop, tip diameter and length, tip cleaning preparation, and whether the experiment was shielded with a Faraday cage. Biasing the tip to different voltages allowed for linear control of the drop charge, and the results were consistent with known electrochemical effects, namely the Volta potential expected between most metals and bulk water ($\approx$ -0.5 V). Introducing insulating plastic materials into the experiment (from the syringe body or tip) imparted large amounts of charge on the drops with apparent stochastic charge evolution. Together these results show that the flow electrification of water is more complex than previously reported, and is driven by various, material-dependent electrostatic processes.
@@ I Introduction
Whether triboelectric charging results from bulk material transfer, electron or ion transfer, from some interfacial adsorbate, or from a combination of all these remains unclear Lacks and Shinbrot (2019). During the last few decades, however, evidence has emerged that water plays a key role in both generating charge separationHarris et al. (2019); Lacks and Shinbrot (2019); Grosjean and Waitukaitis (2023a), stabilizing existing interfacial charge Baytekin et al. (2011), or dissipating charge Méndez Harper et al. (2022, 2024a, 2024b). A salient role of water in triboelectrification is perhaps unsurprising: materials in ambient air generally host layers of absorbed water on their surfaces. Dissociated H+ and OH- ions within these layers may allow for charge exchange during collisions or frictional interactions Gu et al. (2013); Zhang et al. (2015). That water layers modulate charge exchange in systems of solids implies that triboelectrification may also occur at fluid-solid interfaces (or even fluid-fluid interfaces!)Wang (2021); Jin et al. (2022); Armiento et al. (2022). As in solid-solid triboelectrification, the chemical composition of the liquid-solid pair tunes the character of the charging. If the fluid moves relative to the solid, charge separation is known as “flow electrification” Touchard (2001). For viscous fluids, the no-slip assumption implies that the shear forces at the fluid-solid boundary are smaller than those between two frictional solids. However, the contact area is orders of magnitude larger since the liquid conformally coats the solid surface. Even in the absence of flow, there can be a static potential difference between a bulk solid and liquid in contact due to differences in electronic energy states Li et al. (2021).
@@
Flow electrification has been examined for nearly a century, primarily by the petroleum and lubrication engineering industries Klinkenberg and van der Minne (1958); Solomon (1959); ASTM D4865-91 (1991); Harvey et al. (2002); Touchard (2001). Within these contexts, electrostatic charging and discharge pose a significant hazard in dielectric and combustible hydrocarbon fuels Mackeown and Wouk (1942); Bustin and Dukek (1983); Hu et al. (2013). Consider that more than 30 unintended fuel-air mixture ignitions during vehicle refueling were reported in Germany alone between 1992 and 1995. These incidents were ultimately attributed to electrostatic accumulation and subsequent discharge of flowing liquidsvon Pidoll et al. (1997). Charging in dielectric fluids like hydrocarbons is particularly hazardous in filtration systems Gavis and Wagner (1968); Huber and Sonin (1977) and pipe flowBograchev et al. (2012); Abedian and Sonin (1982); Cabaleiro et al. (2008) where the contact area between the liquid and wall is large. For insulating fluids, the Debye layer thickness can extend well beyond the flow boundary layer (more than a few microns). As such, ions in the Debye layer are easily entrained in flow. Additionally, because these fluids have poor conductivities, they do not readily allow for charge recombination. Although this current is often treated phenomenologically, very few studies consider the underlying mechanism of flow electrification Stern (1924); El-Adawy et al. (2011); Touchard et al. (1996); Washabaugh and Zahn (1997).
@@
Flow electrification also occurs in liquids with higher conductivities like water and has recently been proposed as a power source for nanoscopic graphene-based devices Zhang et al. (2021). The Debye length in pure water is $\sim$0.7 nm, orders of magnitude smaller than in dielectric liquids with very few ions. A recent study by Burgo et al. Burgo et al. (2016) presented a triboelectric series for water flowing against materials ranging from air to copper to polytetrafluoroethylene, or PTFE. Both the magnitude and the sign of the flow electrification of water was highly dependent on the composition of the other material. Interestingly, water charged positively against all materials with the exception of air. The fact that air flow can reliably charge water and potentially other liquids is surprising since ambient air contains only $\sim$1 ion/mm3 Méndez Harper et al. (2022). Furthermore, under fair weather conditions, near-surface air contains slightly more positive ions than negative ones Kamra (1982), suggesting that any falling drop would scavenge a net positive charge. In other words, merely collecting ions from the air is insufficient to explain the magnitude and polarity of observed flow electrification between water and air by Burgo and coworkers. Such incongruities, particularly in the context of a ubiquitous material like water, underscore the need for more detailed investigations into contact electrification across fluid-solid and fluid-fluid interfaces.
@@
Here, we report on experiments designed to characterize the charge on individual deionized water drops falling from glass syringes with a variety of needles. The drops, which grow quasi-statically from the tip of the needle, fall into a custom-made Faraday cup, and the mass of each drop is simultaneously measured with a precision balance. In the absence of plastics, we find little or no flow electrification from the ambient environment–that is, the drop charge does not depend on deposition height, metallic tip length, flow rate, or the presence of electrostatic shielding surrounding the experiment. The charge per unit mass on each drop, $\Delta q/\Delta m$, ranged from -5 to -1 pC/g, consistent with the expected Volta potential between typical metals and bulk water Li et al. (2021). Adding salt to the water reduces the overall charge by $\sim$50%. In contrast, introducing common plastic components in the experimental setup (such as a polypropylene syringe body or PTFE syringe tip) can drastically change the charge on each drop. In the case of a PTFE tip, for instance, the absolute charge-to-mass ratio can be as large as 200 pC/g. Furthermore, the charge polarity gained by drops flowing against plastics depends on the material history (e.g. how long fluid has been flowing through a plastic tip) and environmental conditions. Our results suggest that the electrification of individual water drops flowing against metals may be understood from known electrochemical effects. However, interactions with plastics can give rise to more complex charge exchanges whose underlying mechanisms require further clarification.
@@
II Experimental Procedure
@@
To investigate the charging of drops, we employed the setup depicted in Fig. 1a. A programmable syringe pump (Braintree Scientific BS-8000) was fixed vertically above a custom-made Faraday cup (FC) capable of handling liquids. We programmed the pump to dispense individual drops from a luer-lock syringe into the FC at regular intervals. An electrometer (Keithley 6514) connected to the FC using a flexible, low-noise triax cable allowed us to measure charge on falling drops with 10 femtocoulomb resolution. The cup was placed on a precision balance (Ohaus PX623) with milligram resolution. The Faraday cup was connected to an electrometer (Keithley 6514) using a flexible, low-noise triax cable to minimize mechanical strain that could cause inaccurate measurements from the mass balance. This method allowed us to measure the charge-to-mass ratio of each drop, $\Delta q/\Delta m$, and compare our data with that of previous investigations of fluid electrification Burgo et al. (2016); Hendricks Jr (1962); Snarski and Dunn (1991). Lastly, we monitored the ambient conditions using a temperature and humidity sensor (Aosong AHT20). For all experiments, temperature and relative humidity (RH) varied in the range of 20–22C and 20–50%, respectively. The four variables–charge, mass, temperature, and RH–were sampled every $\sim$0.3 s across the duration of an experiment.
@@
Depending on the inner diameter (ID) of a syringe tip (1.35–2.69 mm) and the flow rate of the syringe pump ($\leq$350 $\mu$L/min), drops were produced with a period of 5–10 s, and diameters of 2.2–4.7 mm. Drops were allowed to fall 10–40 cm in air before landing in the FC. For some experiments, we minimized the influence of spurious electric fields by enclosing the entire experimental setup in a grounded Faraday cage made from aluminum wire mesh. Lastly, any metallic parts of the syringe (e.g. the tip or the metallic base of the syringe luer-lock system) were grounded to the outer shell of the FC using a small-gauge, coiled, solid-core copper wire. The solid core coil minimized any mechanical coupling between the syringe and the FC that would cause errors in mass measurements.
@@
The charge $q$ on a drop entering the FC is registered as a step change in the voltage $V$ across the known feedback capacitor $C$ of the electrometer, $V=q/C$. Similarly, the addition of a drop into the FC causes a stepwise jump in the mass reading. Typical recordings of charge (red curve) and mass (blue curve) for two consecutive drops are shown in Fig. 1b. A sudden change in the balance reading indicates an impinging drop. We used the center point on the plateau, $m_{i}$, between two consecutive impinging drops as the average added mass on the balance. Finally, we estimated the mass of an individual drop $\Delta m$ by computing the difference in scale readings between two neighboring plateaus $\Delta m=m_{i}-m_{i-1}$. We used a similar procedure to extract the charge on a falling drop: a sudden change in the electrometer output corresponds to a drop entering the FC. However, electrometers of this sensitivity are prone to drift across the measurement timescales. Thus, we calculated the charge on a drop ($\Delta q$) by subtracting the electrometer reading immediately before and after the arrival of the drop (as detected by the mass change). A zoomed-out time series showing the mass and charge evolution for 41 consecutive drops is shown in Fig. 1c.
@@
Most experiments were conducted with deionized, ultra pure, filtered water (ELGA Veolia Purelab Chorus 1 Reservoir). The water was handled in clean glass containers before being deposited into the testing syringe. However, we also performed experiments with salt water, glycerol, and Fluorinert FC-70. Pure NaCl and glycerol were obtained from Sigma Aldrich, and the Fluorinert was obtained from 3M. To explore the dependence of solid surface properties on the electrification of fluids, we dispensed individual drops from syringe bodies and tips of varied materials. Most experiments used an all-glass, 50 mL capacity syringe (Tomopal). The syringe was cleaned with pure ethanol, rinsed with deionized water, and then dried in an oven at 50C prior to use. A stainless steel luer-lock was mounted at the base of the glass syringe and was connected to various grounded 304 stainless steel or custom nickel-plated stainless steel syringe tips. For nearly all experiments, the tips were cleaned with ethanol and rinsed with water prior to use, but we also tried two other methods of cleaning syringe tips: rinsing them with Neutrad solution and exposing them to oxygen plasma with a custom-built oxygen plasma oven Mouat et al. (2020); Pye et al. (2018). In addition to glass and metals, we conducted experiments with plastic components since these are commonly used in fluid electrification experiments Burgo et al. (2016) and general fluid transport applications. In one set of experiments, we used standard 20 mL polypropylene syringes (Henke-Ject) connected to standard luer-lock syringe tips composed of a polypropylene luer base and a 304 stainless steel, flat-ended cylindrical tube (1.70 mm ID). In another set of experiments, we used the glass syringes, but all-plastic PTFE syringe tips (1.35 mm ID). All metal and plastic syringe tips were obtained from Vita Needle.
III Results and Discussion
@@
Figure 2: Probability distribution functions (PDF) for $\Delta q/\Delta m$ under different conditions using a glass syringe body connected to a 5 cm long, 2.39 mm ID syringe tip. The flow rate was 350 $\mu$L/min, and data was recorded at 44–45% RH. Each distribution is computed over $n\geq$ 410 data points. Each panel represents a different deposition height ($h$), and also includes data taken with a Faraday cage enclosing the syringe pump, cup, and balance assembly (Fig. 1a). The solid lines indicate the mean of each distribution, and the variance in the data primarily comes from the variance in $\Delta q$ since the drop mass was fairly uniform. For completeness, for the $h=$ 10 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.11 $\pm$ 0.02 pC, $\langle\Delta q_{\mathrm{cage}}\rangle=$0.13 $\pm$ 0.01 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.037 $\pm$ 0.006 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.042 $\pm$ 0.005 g. For the $h=$ 20 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.13 $\pm$ 0.04 pC, $\langle\Delta q_{\mathrm{cage}}\rangle=$0.14 $\pm$ 0.02 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.042 $\pm$ 0.009 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.040 $\pm$ 0.004 g. For the $h=$ 30 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.15 $\pm$ 0.04 pC, $\langle\Delta q_{\mathrm{cage}}\rangle=$0.15 $\pm$ 0.03 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.047 $\pm$ 0.005 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.046 $\pm$ 0.003 g. Lastly, for the $h=$ 40 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.14 $\pm$ 0.04 pC, $\langle\Delta q_{cage}\rangle=$0.11 $\pm$ 0.02 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.041 $\pm$ 0.005 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.041 $\pm$ 0.003 g.
@@
The most salient results of our experiments with all glass and metal syringes are illustrated in Fig. 2a-d, which shows histograms of $\Delta q/\Delta m$ for many experimental conditions. We found little to no dependence of $\Delta q/\Delta m$ on the drop deposition height ($h$) or the presence of a Faraday cage shielding the experimental apparatus. The distributions were assembled from multiple independent experiments using 2 mL of water and were completed over multiple days. The average value of $\Delta q/\Delta m$ for each distribution is shown as a vertical line. Most drops fall within the range of -2.0 pC/g to -4.5 pC/g (5th and 95th percentile, respectively). The addition of a protective Faraday cage around the entire setup made a noticeable difference for $h=$ 40 cm, but the shift of the distribution mean was smaller than the standard deviation. Taken together, these results results suggest that the characteristic charge on each drop is determined by the materials in contact and its history, and not its interaction with air as it falls from the syringe needle.
As noted above, we cleaned the syringe needles using three different preparation methods: rinsing them with pure ethanol followed by DI water, soaking them in a Neutrad ultrasonic bath then rinsing with DI water, and exposing them to an oxygen plasma for 30 s. The effects of different cleaning procedures are summarized in Fig. 3. The histograms of $\Delta q/\Delta m$ suggest that the charge gained by water droplets in contact with metal are generally independent of cleaning procedure. We note, however, that the oxygen plasma cleaning did decrease average drop mass, presumably because it changed the contact angle of the water and stainless steel contact line. This affects the formation of the drop near the attachment to the syringe tip, and ultimately the force balance that determines the maximum weight of the drop that can be supported.
@@
The Volta potential is often measured with scanning Kelvin probe force microscopy Ma et al. (2021); Guo et al. (2012); Örnek et al. (2019), and is the difference between the potential of zero charge, $U_{pzc}$, and the work function of the metal, $\Phi$:
@@
The potential of zero charge is analogous to the work function when the metal is in contact with a solution instead of vacuum. The Volta potential can be thought of as the bulk potential of a solution (relative to vacuum) when in contact with a metal surface. From theory, $\Delta\psi$ has two contributions: a reorientation of water molecules at the metal surface, $g^{\text{solv}}(\text{dip})_{0}$, and a redistribution of surface metal electrons, $-\delta\chi^{M}_{0}$ Trasatti and Lust (1999); Mohandas et al. (2024). To further investigate the possibility that $V_{D}$ is a natural electrochemical bias developed between the metal interface and the bulk water, we used a power supply to bias the voltage of our metallic syringe tips (both 304 stainless and nickel-plated stainless steel) during drop deposition into the FC. The results are illustrated in Fig. 4. The voltage of the syringe tip, $V_{\text{tip}}$, was measured relative to the outer grounded shell of the FC, and $V_{D}$ was calculated from Eq. 1.

### v2

# The electrostatic charge on exuded liquid drops

@@ Abstract
Fluid triboelectrification, also known as flow electrification, remains an under-explored yet ubiquitous phenomenon with potential applications from material science to planetary evolution. Building upon previous efforts to position water within the triboelectric series, we investigate the charge on individual, millimetric water drops falling through air. Our experiments measured the charge and mass of each drop using a Faraday cup mounted on a mass balance, and connected to an electrometer. For pure water in a glass syringe with a grounded metal tip, we find the charge per drop ($\Delta q/\Delta m$) was approximately -5 pC/g to -1 pC/g. This was independent of the release height of the drop, tip diameter and length, tip cleaning preparation, and whether the experiment was shielded with a Faraday cage. Biasing the tip to different voltages allowed for linear control of the drop charge, and the results were consistent with known electrochemical effects, namely the Volta potential expected between most metals and bulk water ($\approx$ -0.5 V). Introducing insulating plastic materials into the experiment (from the syringe body or tip) imparted large amounts of charge on the drops with systematic charge evolution. Together these results show that the flow electrification of water is more complex than previously reported, and is driven by material-dependent electrostatic processes.
@@ I Introduction
Whether triboelectric charging results from bulk material transfer, electron or ion transfer, some interfacial adsorbate, or from a combination of all these remains unclear Lacks and Shinbrot (2019). During the last few decades, however, evidence has emerged that water plays a key role in generating charge separationHarris et al. (2019); Lacks and Shinbrot (2019); Grosjean and Waitukaitis (2023a), stabilizing existing interfacial charge Baytekin et al. (2011), and dissipating charge Méndez Harper et al. (2022, 2024a, 2024b). A salient role of water in triboelectrification is perhaps unsurprising: materials in ambient air typically host layers of adsorbed water on their surfaces, unless they are particularly hydrophobic. Dissociated H+ and OH- ions within these layers may allow for charge exchange during collisions or frictional interactions Gu et al. (2013); Zhang et al. (2015). That water layers modulate charge exchange in systems of solids implies that triboelectrification may also occur at fluid-solid interfaces (or even fluid-fluid interfaces!)Wang (2021); Jin et al. (2022); Armiento et al. (2022). As in solid-solid triboelectrification, the chemical composition of the liquid-solid pair tunes the character of the charging. If the fluid moves relative to the solid, charge separation is known as “flow electrification” Touchard (2001). For viscous fluids, the no-slip assumption implies that the shear forces at the fluid-solid boundary are smaller than those between two frictional solids. However, the contact area is orders of magnitude larger than expected for solid-solid frictional contacts Rubinstein et al. (2006) since the liquid coats the entirety of the solid surface. Even in the absence of flow, there can be a static potential difference between a bulk solid and liquid in contact due to differences in electronic energy states Li et al. (2021).
@@
Flow electrification has been examined for more than a century, primarily by the petroleum and lubrication engineering industries Klinkenberg and van der Minne (1958); Solomon (1959); ASTM D4865-91 (1991); Harvey et al. (2002); Touchard (2001). Within these contexts, electrostatic charging and discharge pose a significant hazard in dielectric and combustible hydrocarbon fuels Mackeown and Wouk (1942); Bustin and Dukek (1983); Hu et al. (2013). Consider that more than 30 unintended fuel-air mixture ignitions during vehicle refueling were reported in Germany alone between 1992 and 1995. These incidents were ultimately attributed to electrostatic accumulation and subsequent discharge of flowing liquidsvon Pidoll et al. (1997). Charging in dielectric fluids like hydrocarbons is particularly hazardous in filtration systems Gavis and Wagner (1968); Huber and Sonin (1977) and pipe flowBograchev et al. (2012); Abedian and Sonin (1982); Cabaleiro et al. (2008) where the contact area between the liquid and wall is large. For insulating fluids, the Debye layer thickness can extend well beyond the flow boundary layer (more than a few microns). As such, ions in the Debye layer are easily entrained in flow. Additionally, because these fluids have poor conductivities, they do not readily allow for charge recombination. Although this current is often treated phenomenologically, very few studies consider the underlying mechanism of flow electrification Stern (1924); El-Adawy et al. (2011); Touchard et al. (1996); Washabaugh and Zahn (1997).
Flow electrification also occurs in liquids with higher conductivities like water. In 1892, inspired by observations of negative charge developed in air around waterfalls across the Alps, Lenard conducted a series of laboratory experiments to study the charging behavior of falling water drops or streams in ambient air Lenard (1892). These pioneering experiments showed that the water dispensed from a reservoir collects charge while the surrounding air develops an oppositely-charged potential. Lenard noted variations in this behavior depending on the cleanliness and source of the dispensed water, among other material properties used in the experiments. The Debye length in pure water is $\sim$0.7 nm, orders of magnitude smaller than in dielectric liquids with very few ions. This has led to proposals to use water as a power source for nanoscopic graphene-based devices Zhang et al. (2021). Another recent study by Burgo et al. presented a triboelectric series for water flowing against materials ranging from air to copper to polytetrafluoroethylene, or PTFE. Both the magnitude and the sign of the flow electrification of water was highly dependent on the composition of the other material Burgo et al. (2016).
@@
Interestingly, Burgo et al. showed that water charged positively against all materials with the exception of air. The fact that air flow can reliably charge water and potentially other liquids is surprising since ambient air contains only $\sim$1 ion/mm3 Méndez Harper et al. (2022). Furthermore, under fair weather conditions, near-surface air contains slightly more positive ions than negative ones Kamra (1982), suggesting that any falling drop would scavenge a net positive charge. In other words, merely collecting ions from the air is insufficient to explain the magnitude and polarity of observed flow electrification between water and air by Burgo and coworkers. Such incongruities, particularly in the context of a ubiquitous material like water, underscore the need for more detailed investigations into contact electrification across fluid-solid and fluid-fluid interfaces.
@@
Here, we report on experiments designed to characterize the charge on individual deionized water drops falling from glass syringes with a variety of needles. The drops, which grow quasi-statically from the tip of the needle, fall into a custom-made Faraday cup, and the mass of each drop is simultaneously measured with a precision balance. In the absence of plastics, we find little or no flow electrification from the ambient environment–that is, the drop charge does not depend on deposition height, metallic tip length, flow rate, or the presence of electrostatic shielding surrounding the experiment. The charge per unit mass on each drop, $\Delta q/\Delta m$, ranged from -5 to -1 pC/g, consistent with the expected Volta potential between typical metals and bulk water Li et al. (2021). Adding salt to the water reduces the overall charge by $\sim$50%. In contrast, introducing common plastic components in the experimental setup (such as a polypropylene syringe body or PTFE syringe tip) can drastically change the charge on each drop. In the case of a PTFE tip, for instance, the absolute charge-to-mass ratio can be as large as 120 pC/g. Furthermore, the charge polarity gained by drops flowing against plastics depends on the material history (e.g. how long fluid has been flowing through a plastic tip) and environmental conditions. Our results suggest that the electrification of individual water drops flowing against metals may be understood from known electrochemical effects. However, interactions with plastics can give rise to more complex charge exchanges whose underlying mechanisms require further clarification.
@@
Experimental Procedure
To investigate the charging of drops, we employed the setup depicted in Fig. 1a. A programmable syringe pump (Braintree Scientific BS-8000) was fixed vertically above a custom-made Faraday cup (FC) capable of handling liquids. We programmed the pump to dispense individual drops from a luer-lock syringe into the FC at regular intervals. The FC consisted of an inner sensing cup nested within a grounded shielding cylinder separated by a PTFE spacers. The inner cup was electrically connected to the center pin a triax bulkhead connector, while the outer housing was grounded. The inner cup had a diameter of 4.57 cm and a depth of 3.05 cm, yielding a total internal volume of 50 mL. Individual FC components were ultrasonically cleaned, dried in an oven, and allowed to cool before assembly and employment in our experimental setup. An electrometer (Keithley 6514) connected to the FC using a flexible, low-noise triax cable (operating in the ¡20 nC range) allowed us to measure charge on falling drops with a resolution of 0.01 pC.
@@
To measure the mass of individual drops, the FC was placed on a precision balance (Ohaus PX623) with milligram resolution. Thus, we were able to measure the charge-to-mass ratio of each drop, $\Delta q/\Delta m$, and compare our data with that of previous investigations of fluid electrification Burgo et al. (2016); Hendricks Jr (1962); Snarski and Dunn (1991). Lastly, we monitored the ambient conditions using a temperature and humidity sensor (Aosong AHT20). For all experiments, temperature and relative humidity (RH) varied in the range of 20–22C and 20–50%, respectively. The four variables–charge, mass, temperature, and RH–were sampled every $\sim$0.3 s across the duration of an experiment.
@@
Depending on the inner diameter (ID) of a syringe tip (1.35–2.69 mm) and the flow rate of the syringe pump, drops were produced with a period of 5–10 s and radius of 1.85–2.35 mm. For most trials, we employed a flow rate of 350 $\mu$L/min, but also tested a range of flow rates (35–750 $\mu$L/min) were individual drops could be resolved by the mass balance. These results are shown in Fig. 1e-f, and show little variation in drop charge and mass with the flow rate. Drops were allowed to fall 10–40 cm in air before landing in the FC. For some experiments, we minimized the influence of spurious electric fields by enclosing the entire experimental setup in a grounded Faraday cage made from aluminum wire mesh. Lastly, any metallic parts of the syringe (e.g. the tip or the metallic base of the syringe luer-lock system) were grounded to the outer shell of the FC using a small-gauge, coiled, solid-core copper wire. The solid core coil minimized any mechanical coupling between the syringe and the FC that would cause errors in mass measurements.
@@
The charge $q$ on a drop entering the FC is registered as a step change in the voltage $V$ across the known feedback capacitor $C$ of the electrometer, $V=q/C$. Similarly, the addition of a drop into the FC causes a stepwise jump in the mass reading. Typical recordings of charge (red curve) and mass (blue curve) for two consecutive drops are shown in Fig. 1b. A sudden change in the balance reading indicates an impinging drop. To allow for stabilization after impact, we measured the center point on the plateau, $m_{i}$, between two consecutive impinging drops, and estimated the mass of an individual drop $\Delta m$ by computing the difference in readings between two neighboring plateaus $\Delta m=m_{i}-m_{i-1}$. We used a similar procedure to extract the charge on a falling drop: a sudden change in the electrometer output corresponds to a drop entering the FC. However, electrometers of this sensitivity are prone to drift across the measurement timescales. Thus, we calculated the charge on a drop ($\Delta q$) by subtracting the electrometer reading immediately before and after the arrival of the drop (as detected by the mass change). A zoomed-out time series showing the mass and charge evolution for 41 consecutive drops is shown in Fig. 1c.
@@
Most experiments were conducted with deionized, ultra pure, filtered water (ELGA Veolia Purelab Chorus 1 Reservoir). The water was handled in clean glass containers before being deposited into the testing syringe. However, we also performed experiments with salt water, glycerol, and Fluorinert FC-70. Pure NaCl and glycerol were obtained from Sigma Aldrich, and the Fluorinert was obtained from 3M. To explore the dependence of solid surface properties on the electrification of fluids, we dispensed individual drops from syringe bodies and tips of varied materials. Most experiments used an all-glass, 50 mL capacity syringe with a stainless steel luer-lock mounted at the base (Tomopal). The syringe was cleaned with pure ethanol, rinsed with deionized water, and then dried in an oven at 50C prior to use. The glass syringe was connected to various grounded 304 stainless steel or custom nickel-plated stainless steel syringe tips. For nearly all experiments, the tips were cleaned with ethanol and rinsed with water prior to use, but we also tried two other methods of cleaning syringe tips: rinsing them with Neutrad solution and exposing them to oxygen plasma with a custom-built oxygen plasma oven Mouat et al. (2020); Pye et al. (2018). In addition to glass and metals, we conducted experiments with plastic components since these are commonly used in fluid electrification experiments Burgo et al. (2016) and broad fluid transport applications. In one set of experiments, we used 20 mL polypropylene syringes (Henke-Ject) connected to standard luer-lock syringe tips composed of a polypropylene luer base and a 304 stainless steel, flat-ended cylindrical tube (1.70 mm ID). In another set of experiments, we used the glass syringes, but all-plastic PTFE syringe tips (1.35 mm ID). All metal and plastic syringe tips were obtained from Vita Needle.
Results and Discussion
@@
Figure 2: Probability distribution functions (PDF) for $\Delta q/\Delta m$ using a glass syringe body connected to a 5 cm long, 2.39 mm ID syringe tip. The flow rate was 350 $\mu$L/min, and data was recorded at 44–45% RH. Each distribution is computed over $n\geq$ 410 data points. Each panel (a-d) represents a different deposition height ($h$), and also includes data taken with a Faraday cage enclosing the syringe pump, cup, and balance assembly (Fig. 1a). The solid lines indicate the mean of each distribution, and the variance in the data primarily comes from the variance in $\Delta q$ since the drop mass was fairly uniform. For completeness, for the $h=$ 10 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.11 $\pm$ 0.02 pC, $\langle\Delta q_{\mathrm{cage}}\rangle=$0.13 $\pm$ 0.01 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.037 $\pm$ 0.006 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.042 $\pm$ 0.005 g. For the $h=$ 20 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.13 $\pm$ 0.04 pC, $\langle\Delta q_{\mathrm{cage}}\rangle=$0.14 $\pm$ 0.02 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.042 $\pm$ 0.009 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.040 $\pm$ 0.004 g. For the $h=$ 30 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.15 $\pm$ 0.04 pC, $\langle\Delta q_{\mathrm{cage}}\rangle=$0.15 $\pm$ 0.03 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.047 $\pm$ 0.005 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.046 $\pm$ 0.003 g. Lastly, for the $h=$ 40 cm distributions, we found $\langle\Delta q_{\mathrm{air}}\rangle=$0.14 $\pm$ 0.04 pC, $\langle\Delta q_{cage}\rangle=$0.11 $\pm$ 0.02 pC, $\langle\Delta m_{\mathrm{air}}\rangle=$ 0.041 $\pm$ 0.005 g, and $\langle\Delta m_{\mathrm{cage}}\rangle=$ 0.041 $\pm$ 0.003 g. (e) Plot of $\langle\Delta q\rangle$ against flow rate. (f) Plot of $\langle\Delta m\rangle$ against flow rate. Error bars were produced from the standard deviation of per-trial means at each flow rate, and all trials were completed at $h=$ 10 cm.
@@
The aggregated results of our experiments with all glass and metal syringes are illustrated in Fig. 2a-d, which shows histograms of $\Delta q/\Delta m$ for many experimental conditions. We found little to no dependence of $\Delta q/\Delta m$ on the drop deposition height ($h$) or the presence of a Faraday cage shielding the experimental apparatus. The distributions were assembled from multiple independent experiments using 2 mL of water and were completed over multiple days. The average value of $\Delta q/\Delta m$ for each distribution is shown as a vertical line. Most drops fall within the range of -2.0 pC/g to -4.5 pC/g (5th and 95th percentile, respectively). The addition of a protective Faraday cage around the entire setup made a noticeable difference for $h=$ 40 cm, but the shift of the distribution mean was smaller than the standard deviation. Taken together, these results suggest that the characteristic charge on each drop is determined by the materials in contact and its history, and not its interaction with air as it falls from the syringe needle.
As noted above, we cleaned the syringe needles using three different preparation methods: rinsing them with pure ethanol followed by DI water, soaking them in a Neutrad ultrasonic bath then rinsing with DI water, and exposing them to an oxygen plasma for 30 s. The effects of different cleaning procedures are summarized in Fig. 3. The histograms of $\Delta q/\Delta m$ suggest that the charge gained by water droplets in contact with metal are independent of cleaning procedure. We note, however, that the oxygen plasma cleaning did decrease average drop mass, presumably because it changed the contact angle of the water and stainless steel contact line. This affects the formation of the drop near the attachment to the syringe tip, and ultimately the force balance that determines the maximum weight of the drop that can be supported.
@@
The Volta potential is often measured with scanning Kelvin probe force microscopy Ma et al. (2021); Guo et al. (2012); Örnek et al. (2019), and is defined as the difference between the potential of zero charge, $U_{pzc}$, and the work function of the metal, $\Phi$:
@@
The potential of zero charge is analogous to the work function when the metal is in contact with a solution instead of vacuum. The Volta potential can be thought of as the bulk potential of a solution (relative to vacuum) when in contact with a metal surface. From theory, $\Delta\psi$ has two contributions: a reorientation of water molecules at the metal surface, $g^{\text{solv}}(\text{dip})_{0}$, and a redistribution of surface metal electrons, $-\delta\chi^{M}_{0}$ Trasatti and Lust (1999); Mohandas et al. (2024). To further investigate the possibility that $V_{D}$ is a natural electrochemical bias developed between the metal interface and the bulk water, we used a power supply to bias the voltage of our metallic syringe tips (both bare and nickel-plated 304 stainless steel) during drop deposition into the FC. The results are illustrated in Fig. 4. The voltage of the syringe tip, $V_{\text{tip}}$, was measured relative to the outer grounded shell of the FC, and $V_{D}$ was calculated from Eq. 1.
Figure 5: Diagram of a potential charging mechanism for pure water in contact with a conducting metal surface. Paired water molecules self-ionize into hydroxide (OH-) and hydronium (H3O+) ions. A small fraction (less than 1 ppm) of H3O+ ions source an electron (e-) from the metal, and ultimately produce neutral hydrogen gas (H). This leads to a net charge imbalance where the hydroxide ions preferentially accumulate at the air-water interface of the drop. This process continues until the drop has reached its equilibrium Volta potential, $\Delta\psi$, between the water and metal. The neutral hydrogen will eventually form diatomic hydrogen (H2) and escape the drop into ambient air.